At 570. mm Hg and 25 °C, a gas sample has a volume of 2270 mL. What is the final pressure (in mmHg) at a volume of 125 mL and a temperature of 17.5 °C?
Answer #1

Answer:

Since P1V1/T1=P2V2/T2

Where P=pressure, V=volume and T=temperature

Given P1=570 mmH V1=2270 mL, T1=25°C=25+273 K=298 K.

P2=?, V2=125 mL, and T2=17.5°C=17.5+273 K=290.5 K.

P2=P1V1T2/T1V2

P2=(570 mmH x 2270 mL x 290.5 K)/(298 K x 125 mL)

P2=10090.68 mm Hg.

Please let me know if you have any doubt. Thanks